A store reduces the price of an MP3 player by 20% and then reduces that price by 15%. If the final price of the MP3 player is $170, what was its original price?

Found 2 solutions by Edwin McCravy, dkppathak:
Answer by Edwin McCravy(20055)

When you reduce an amount by 20% you take 80% of it
When you reduce an amount by 15% you take 85% of it.

0.85(0.80)x = 170
      0.68x = 170
          x = 170/0.68
          x = 250

Answer:  $250

Edwin

Answer by dkppathak(439)

A store reduces the price of an MP3 player by 20% and then reduces that price by 15%. If the final price of the MP3 player is $170, what was its original price?
solution
Let initial cost of MP3 player is $x
as per condition 20% reduction =0.8x
again reduction 15%=0.8x*0.85= 0.68x
0.68x=170
x=170/0.68=$250
cost of player is $250